OpenStudy (anonymous):

Graph the line with slope 4/3 that passes through (3, 1). Refer to Example 2 in page 106 of your textbook for help.

OpenStudy (mathlegend):

Use point slope form.

OpenStudy (mathlegend):

y-y1=m(x-x1) We have x1, y1 and we have a slope 4/3

OpenStudy (mathlegend):

y-1=4/3(x-3)

OpenStudy (anonymous):

y=4/3x-9

OpenStudy (mathlegend):

y-1=4/3x-4 +1 +1 ---------- y = 4/3x - 3

OpenStudy (mathlegend):

@khem-de-economist you did something wrong in your calculation.

OpenStudy (mathlegend):

@yarii do you understand?

OpenStudy (anonymous):

yeah a little i think i got it tho thanks tho

OpenStudy (mathlegend):

:)